OpenStudy (anonymous):

sec^2x-1/tan/2 sin cos

OpenStudy (luigi0210):

what are we doing?

OpenStudy (anonymous):

my question is all about limits kindly help me.. lim sec^2x-1/tan/2sin cos x=0

OpenStudy (luigi0210):

too many slashes >.<

OpenStudy (anonymous):

ya ..

OpenStudy (luigi0210):

Can you use the equation editor thing? My tiny brain cannot comprehend it.

OpenStudy (anonymous):

ok i ll try

OpenStudy (luigi0210):

Me? Pfft I hope you're kidding D:

OpenStudy (anonymous):

\[\lim_{x=0 \rightarrow ?} \frac{ \sec^2x-1 }{ ?tanx}/2 \sin \cos\]

OpenStudy (luigi0210):

You could draw it too if you think you can :P

OpenStudy (anonymous):

looks like that way but 2sin cos is under tanx

OpenStudy (anonymous):